If you mix 60mL of 0.2 M acetic acid with 80mL of 0.4 M acetate, what will be the resulting pH. Then how many total milliliters of 1M NaOH can you add to the solution and still have a good buffer? (that is, within 1 pH unit of pKa).
Answer #1

1)

mmoles of acetic acid = 60 x 0.2 = 12

mmoles of acetate = 80 x0.4 = 32

pH = pKa + log (acetate / acetic acid )

pH = 4.74 + log(12 / 32)

pH = 4.31

2)

pH   = pKa + log (acetate + V / acetic acid - V)

5.74 = 4.74 + log (acetate + V / acetic acid - V)

(acetate + V / acetic acid - V) = 10

12 + V / 32 - V = 10

12 + V = 320 - 10V

11 V = 308

V = 28 mL

volume of NaOH needed = 28 mL